Status Report

NASA Space Station On-Orbit Status 9 Feb 2004

By SpaceRef Editor
February 9, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 9 Feb 2004

All ISS systems continue to function nominally, except as noted previously or below.  Underway: Week 16 of Increment 8. 

FE Alexander Kaleri successfully removed and replaced the Vozdukh’s noisy MR5-15G blower fan.  The carbon dioxide (CO2) scrubber was then again switched to automatic mode, with CO2 partial pressure set at 3.5 mmHg and airflow at 80%.  During the activity, dust stirred up by the crew caused a close-by smoke detector to issue a false fire alarm.  Appropriate procedures recovered and restored all affected systems.   [The loud noise from the Vozdukh CO2 scrubber was heard by the crew first on 2/2.  The source was identified as the MR5 “microcompressor”, which Kaleri replaced today with the onboard spare.]

Using the SSC7 (station support computer #7) laptop in the Airlock (A/L), CDR/SO Michael Foale initiated a maintenance charge/discharge cycle on the two new EMU/spacesuit batteries delivered on Progress 13P, #2045 & #2046, in the A/L’s BSA (battery stowage assembly).  [The charging will take about 24 hrs and will be followed by discharge.  Helmet light and PGT (pistol grip tool) batteries were not charged, and Mike had to verify beforehand that they were not in the BSA.  EMU (extravehicular mobility unit) battery maintenance is performed every 50 days, and consists of fully charging and then discharging the batteries to prolong their useful life.]

Kaleri terminated the “bake-out” cycle on the BMP micropurification system’s channel #2, moding the channel back to Purify.  After yesterday’s termination of regeneration on channel #1, this restores both filter beds to Purification/Absorption mode.  [The regeneration of the air purifier filter beds is repeated every 20 days.  Each bakeout to space vacuum takes about 24 hours.]

The crew continued cargo transfers from the Progress, assisted by a detailed Russian & US transfer/unpack list and the IMS (inventory management system).  Transfers will dominate the crew’s timeline tomorrow and Wednesday.   [U.S. and Russian specialists, supported by Mike’s downlinked video footage, are working on sophisticated planning to alleviate the growing stowage problem, particularly in the U.S. segment (USOS).  Extended stowage planning is also being developed for the 8S docked mission, to better organize the necessary complex crew handover & cargo transfer activities.]

The FE completed another monthly sensor reading of the “Pille-MKS” radiation dosimeter, with its ten sensors (one on stand-by) placed at various locations in the Russian segment (port cabin window, starboard cabin window, ASU toilet facility, control panel, etc.).  Pille dosage values were to be called down to MCC-M or downlinked via Regul comm.  [Last time done: 1/14.]

Mike Foale transferred exercise data files from the TVIS treadmill to the medical equipment computer (MEC) for subsequent downlink and completed the weekly TVIS maintenance, including the regular inspection of the wire ropes for signs of fraying.

Sasha Kaleri broke out and assembled the equipment of the Russian BIO-5 Rasteniya-2/Lada-2 (“Plants-2”) experiment that researches plant growth and development under spaceflight conditions.  The activity was supported by tagup with specialists at TsUP via S-band.   [In the Rasteniya “greenhouse”, experimental seeds of acacia-leaf type pea will be planted between wicks in a root tray, with environmental control powered on.  Regular daily maintenance of the experiment involves monitoring of seedling growth, humidity measurements, moistening of the substrate if necessary, and photo/video recording.]

At the Lab RWS (robotics work station), Mike Foale connected the UOP-DCP (utility outlet panel-to-display & control panel) bypass power cable for the MSS (mobile service system).  Later, he completed a 1-hr. software review of the Dynamic Operational Ubiquitous Graphics (DOUG) application for tomorrow’s DOUG setup and Robotics/SSRMS proficiency training.   [For covering the EVA-9 on 2/26 with its video systems, the SSRMS (space station remote manipulator system) will be “walked off” inchworm-like onto the Lab PDGF (power & data grapple fixture).  DOUG is a sophisticated software program that provides a graphic birdseye-view graphical image of the SSRMS, showing its real-time location and configuration on a PC during its operation.]

The Science Officer used the EHS SSK (environmental health systems/surface sample kit) to perform surface sample swabs in Lab, SM and Node for cultivation.   [SSK sampling is done once per month for the first three months that a module is on orbit and once every three months thereafter.  Bacterial and fungal samples are taken at two locations in each module.  The colony growth on the sampling slides will be analyzed after five days of incubation.]

Foale also employed the MAS (microbial air sampler) kit to collect air samples in Lab, Node and SM for bacterial and fungal analysis.   [As for the SSK, the sampling analysis is done once per month for the first three months, and once every three months thereafter.  Bacterial and fungal air samples are taken at two locations in each module.  The colony growth on the sampling slides will be analyzed after five days of incubation in Petri dishes.]

Mike unpacked the new 13P-delivered scopemeter, which is fully calibrated and replaces the “Pilobolus” as the primary scopemeter for all electric measurement operations.  If required, Foale was to install its NiCd battery pack for a recharge in the SMPA (SM power adapter) battery charger, which is connected with US cables to the Russian power system.  [The SMPA transforms the Russian 28 volts direct current to 12VDC for the battery.  The charging procedure extends over 21 hours.]

The crew did their nominal physical exercise on the TVIS treadmill, RED exerciser and VELO cycle with load trainer.

Mike conducted the regular routine status checkup on autonomously running Inc. 8 payloads in the Lab, while Sasha completed the daily routine maintenance of the SM’s SOZh life support system and prepared the daily IMS delta file for automatic export/import to update the database.

The CDR reported a failed GLA (general luminaire assembly) light fixture in the Lab and requested development of a plan for better utilization of all GLAs remaining onboard, to prevent reduction of lighting conditions in the U.S. segment (USOS).

Specialists at MCC-Houston are initiating new ground-based troubleshooting steps on the Ku-band “forward” link.  The aim is to isolate the problem(s) associated with OCA (orbital communication adapter) not having good two-way communications at all times during good Ku-band passes.   [OCA will be affected no more than 10 minutes at a time for the test period, which is to last approximately the next month).]

In preparation for the Lab science window depressurization activity on Wednesday (2/11), the ground performed thermal characterization on the window, consisting of heating the window in the range of 30.0-37.0 deg C for verifying the window heating characteristics and testing the heater itself.

Digital images of the thin strip-like piece of FOD (foreign object debris) seen floating away from the ISS on 2/5 after the Progress thruster tests have been sent to Moscow for analysis.  Neither U.S. nor Russian specialists have as yet identified the object.  Analysis is continuing.  No systems anomaly that could be associated with the FOD has been observed to date.

Today’s CEO (crew earth observations) targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Inland deltas, SE Sudan (ideal sunglint pass over the vast region of inland deltas of SE Sudan.  The crew could see first the Nile’s inland delta [Sudd swampland]: glint images showed present active streambeds, which are known to change position on the delta surface.  Of greater interests are a series of newly identified, unmapped deltas immediately east of the Nile, banked against the Ethiopian plateau), Bahamas (400mm-lens: Pass over the central part of the island chain; near-nadir views requested), Internal waves, S Mexico (ideal sunglint pass looking right.  In this part of the easternmost Pacific, internal waves have been documented but the cause of the waves is still not known. Unmapped seamounts have been mentioned as a possibility.  [Further south off Panama, internal wave packets are known to be related to the submarine “Cocos Ridge”, an extension of the Galapagos chain of sub sea mountains.]), and Albuquerque, New Mexico (nadir pass).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 2:03pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On.  Vozdukh CO2 scrubber is On (in Automatic Mode).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is off (in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off (repair incomplete).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 25.6; ppO2 (mmHg) — 157.2; ppCO2 (mmHg) — 2.5;
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C) — 20.4.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 23.7.
  • Node:  Pressure (mmHg) — 748.34; temperature (deg C) — 23.7 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 750.88; temperature (deg C) — 25.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 750.98; temperature (deg C) — 28.5; shell heater temp (deg C) — 25.1, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 23.4
  • PMA-2:  Shell heater temp (deg C)  11.9

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed Position (non-suntracking, “night glider”/”sun slicer” drag reduction mode).
  • SM batteries:  Battery #8, formerly known as #7, is failed (to be replaced); battery #5 is in “Cycle” mode; all other batteries (6) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #3 is off (capacity restoration mode, ROM); all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available: 4115 kg (9072 lb) as of 2/5/04  [SM(755) + FGB(2701) + Progress M(0) + Progress M-1(659)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -90 deg, pitch: -9 deg, roll: 1.7 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #1/LEE B, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 8:04am EST [= epoch]):

  • Mean altitude — 367.6 km
  • Apogee — 372.7 km
  • Perigee — 362.6 km
  • Period — 91.9 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0007429
  • Orbits per 24-hr. day — 15.67
  • Mean altitude loss last 24 hours — 80 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 29813

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.